# Tailoring Dynamical Codes for Biased Noise: The X3Z3 Floquet Code

### v1

@@ Abstract
We propose the X3Z3 Floquet code, a type of dynamical code with improved performance under biased noise compared to other Floquet codes. The enhanced performance is attributed to a simplified decoding problem resulting from a persistent symmetry under infinitely biased noise, which suprisingly exists in a code without constant stabilisers. Even if such a symmetry is allowed, we prove that a general dynamical code with two-qubit parity measurements cannot admit one-dimensional decoding graphs, a key feature resulting in the high performance of bias-tailored stabiliser codes. Despite this limitation, we demonstrate through our comprehensive numerical simulations that the symmetry of the X3Z3 Floquet code renders its performance under biased noise far better than several leading Floquet code candidates. Furthermore, to maintain high-performance implementation in hardware without native two-qubit parity measurements, we introduce ancilla-assisted bias-preserving parity measurement circuits. Our work establishes the X3Z3 code as a prime quantum error-correcting code candidate, particularly for devices with reduced connectivity, such as the honeycomb and heavy-hexagonal architectures.
@@ I Introduction
Quantum error correction (QEC) [1, 2, 3, 4] should be understood as occurring both in space and time [5]. Taking advantage of the temporal dimension, Floquet codes [6, 7, 8, 9, 10, 11, 12, 13, 14, 15], or more generally dynamical codes [16, 17], form a large class of error-correcting codes, which can achieve competitive fault-tolerant performance while reducing the weights of check measurements performed during error-correction [6, 7, 9, 18]. Several of these codes [8, 9, 10, 11] also benefit from being defined on a lattice with sparser connectivity than that for the surface code [19, 20]: each qubit is only connected to three other qubits. In architectures where two-qubit parity check measurements are native, Floquet codes could achieve higher thresholds [21, 7] than the surface code. Without requiring additional connectivity, this code can be deformed around defective components due to highly noisy qubits or gates [18, 22]. Moreover, dynamical codes allow for implementations of arbitrary Clifford, and even some non-Clifford gates, through low-weight parity check measurements [16].
@@
Although Floquet codes (in particular, the honeycomb code) have been studied under various noise models [9, 23], there have not been any Floquet codes that are specifically tailored for an improved performance under biased noise. A biased noise model is one in which a specific type of error, for example, phase errors, occurs more frequently than other errors, such as bit flip errors. This biased noise is typical to most quantum platforms, for example, superconducting fluxonium qubits [24], bosonic “cat” qubits [25, 26], neutral atoms [27], quantum-dot spin qubits [28, 29] and Majorana qubits [30].
@@
For enhanced performance, quantum-error-correcting codes need to be designed such that they possess symmetries that can be utilised to simplify the error-syndrome decoding problem given the noise structure [31, 32, 33, 34, 35, 36, 37, 38, 39]. While there have been several proposals on bias-tailored static codes [32, 33, 35, 36, 40, 34, 41, 35, 37, 38, 42, 43, 39], designing Floquet codes for high performance under biased noise is still an open problem. Owing to the experimental relevance of biased noise and given the ease of implementation of Floquet codes which require only two-qubit parity measurements, it is therefore imperative to tailor Floquet codes for biased noise and study how the performance of such dynamical codes can be improved.
@@
In this paper, we present the X3Z3 bias-tailored Floquet code, a Clifford-deformed [36] version of the Calderbank–Shor–Steane (CSS) Floquet code [10, 11]. Despite not having a fixed stabiliser group (as static codes have), the X3Z3 Floquet code still possesses a symmetry under infinitely biased noise that allows for a simplified decoding in biased noise regimes. We perform an in-depth study of this code, along with the CSS Floquet code [10, 11], and two types of honeycomb codes: one proposed by Hastings and Haah [6] and the other by Gidney et al. [7]. We simulate all codes under biased-noise models, and find that the X3Z3 code has the best performance. Using a matching decoder, we find that, as the noise changes from fully depolarising to pure dephasing, the X3Z3 Floquet code threshold increases from $1.13\%$ to $3.09\%$ under a code-capacity noise model and increases from $0.76\%$ to $1.07\%$ under a circuit-level noise model mimicking hardware with noisy direct entangling measurements. Furthermore, we show that its sub-threshold performance is also substantially better under biased noise than other Floquet codes.
@@
Compared to its static counterparts, the X3Z3 Floquet code has an advantage that it can be realised using only two-qubit parity check measurements. This makes it particularly suitable for devices with constrained connectivity, such as the honeycomb and heavy-hexagonal lattice (currently IBM’s preferred superconducting-qubit architecture) [44, 45]. Moreover, we demonstrate that the two-qubit parity measurements of the Floquet code can be performed in a bias-preserving way even in hardware without direct entangling measurements, thus enabling high performance implementation in such devices.
@@
The paper is laid out as follows. In Sec. II, we review the basics of Floquet codes together with two commonly studied examples: honeycomb and CSS Floquet codes. Readers who are already familiar with Floquet codes can skip directly to Sec. III where we discuss our X3Z3 Floquet code which is obtained by bias-tailoring the CSS Floquet code. Crucially, we show that there exists a persistent symmetry in the code’s error syndrome under the pure dephasing noise model that allows for simplified decoding. Subsequently, we introduce ancilla-assisted bias-preserving parity measurement circuits that allow for high-performance in devices without native entangling measurements. In Sec. IV, we present our simulation methods and noise models. We then provide our simulation results for all of the four Floquet codes studied in Sec. V. To explain the code performance, in Sec. VI we prove that a class of dynamical codes on the honeycomb lattice has high-connectivity decoding graphs. We also prove that dynamical codes with only two-qubit parity check measurements cannot have repetition-code-like decoding graphs, as is the case with static codes. Finally, we conclude and present future research directions in Sec. VII. In the Appendices, we give a more detailed review of the basics of honeycomb and CSS Floquet codes, and present the details of our noise models, numerical simulations, and theorems.
@@ II.1 Dynamical and Floquet codes
Figure 1: (a) CSS and (b,c) X3Z3 Floquet codes. (a,b) Left: Hexagonal lattice with qubits sitting on vertices and opposite boundaries identified. Plaquettes are assigned one of three colours (red, green, blue) and one of two types (X- or Z-type for CSS and A- or B-type for X3Z3). (a) Right: CSS Floquet code measurement schedule. (b) The X3Z3 Floquet code can be obtained from the CSS code via Hadamard gates applied to shaded domains shown (left). Right: Plaquette and edge (check) operators are A- or B-type, depending on their support on shaded/unshaded domains. (c) The X3Z3 Floquet code measurement schedule. Arrows indicate the type and colour of the edge operator measured at each step, where the edges just measured (members of the ISG) are highlighted in the lattice. Uncoloured plaquettes host only a single type of stabiliser, either A or B, indicated by the letters in the plaquettes, while coloured plaquettes host both A- and B-type stabilisers for the ISG. One set of anti-commuting logical operators is shown by yellow and light blue strings, where their $X$ ($Z$) support on qubits is highlighted in red (blue). The other set (not shown) is similar to the set shown but offset by three measurement subrounds and with the $X\leftrightarrow Z$ interchange of the qubit supports’ bases.
@@
We begin by defining dynamical and Floquet codes. In this paper, we consider the Floquet codes to be defined on the lattice of a two-dimensional colour code, which is trivalent and three-colourable. A trivalent lattice has each vertex incident to three edges, and a three-colourable lattice has every face assigned one of three colours in such a way that there are no two adjacent faces of the same colour. Throughout this paper, we will use the honeycomb lattice as an example of such a lattice (see Fig. 1).
@@
We define qubits as residing on the vertices of the lattice and error-correction will be performed by measuring two-qubit operators defined on edges of the lattice (i.e., acting on qubits incident to a given edge). Edges may be assigned a colour: each edge is coloured the same as the plaquettes it connects. We perform error-correction by measuring edge operators in a particular sequence. After any given round of edge measurements, the system will be in the $+1$-eigenstate of the operators in an “instantaneous stabiliser group” (ISG), which will change at each time step. The ISG at time step $t+1$ is defined as:
@@
$\begin{split}\text{ISG}_{t+1}=\langle S,\pm M\;|\;M\in\mathcal{M}_{t},\;S\in\text{ISG}_{t},\\ \text{such that }[S,M]=0,\;\forall M\in\mathcal{M}_{t}\rangle.\end{split}$ (1)
@@
In the above, $\mathcal{M}_{t}$ is the set of “check” measurements performed at time step $t$. The ISGt+1 also includes “plaquette stabiliser operators” $S\in\text{ISG}_{t}$ which commute with all $M\in\mathcal{M}_{t}$. The check measurement operators are chosen in such a way that those check operators at time $t$ that have overlapping qubit supports with check operators at $t+1$ anti-commute. For Floquet codes, the measurement sequence is periodic, such that $\mathcal{M}_{t+T}=\mathcal{M}_{t}$ for some integer $T$. For such a code, we will be performing quantum memory experiments with $mT$ time steps, for some integer $m$. We will refer to $m$ as the number of “QEC rounds” in the experiment, while we will refer to $mT$ as the number of “measurement subrounds” in the experiment.
@@
The logical operators at time $t$ are given by $\mathcal{C}(\text{ISG}_{t})/\text{ISG}_{t}$, where $\mathcal{C}(\text{ISG}_{t})$ is the centraliser of ISGt, i.e., the group of Pauli operators commuting with all $S\in\text{ISG}_{t}$. A (potentially trivial) logical operator “representative” is some member of $\mathcal{C}(\text{ISG}_{t})$. Each nontrivial logical operator representative for the codes considered will be a string-like Pauli operator at each time step (see Fig. 1). To avoid anti-commuting with the next-subround edge measurements, certain check measurement results along a logical operator’s path have to multiplied into that logical operator. Hence, the logical operators will evolve from one time step to the next.
@@
We can detect errors if we can find sets of measurements, called detectors, that always multiply to the value $+1$ in the absence of noise, thus registering no error. Over some number of QEC rounds we will have extracted several detector outcomes. A detector (or decoding) hypergraph is formed by first defining a node for each (independent) detector in the code’s history. Subsequently, for each potential fault (e.g., Pauli or measurement errors) that might have occured, a (hyper)edge is drawn between the detectors whose signs are flipped by this fault. Each (hyper)edge is assigned a weight based on the probability of the corresponding error occurring [46]. The codes we will be examining are amenable to minimum-weight perfect matching decoding [47], upon decomposing hyperedges into edges. Given a “syndrome” (a set of detectors whose measurements return $-1$ rather than $+1$), the decoder attempts to pair up the triggered detectors to determine a shortest-path correction operation. The decoder succeeds if the error combined with the correction is a trivial logical operator.
@@ II.2 Honeycomb codes
We define one stabiliser operator for each plaquette, such that blue plaquettes host $Z^{\otimes 6}$ operators, red plaquettes host $X^{\otimes 6}$ operators and green plaquettes host $Y^{\otimes 6}$ operators. These plaquette operators are inferred from edge measurements in two consecutive rounds. Detectors are formed from consecutive plaquette operator measurements.
@@
The second honeycomb code variant, which we call the XYZ2 honeycomb code, is due to Hastings and Haah [6]. It differs from the P6 code by single-qubit Clifford rotations acting on the qubits. While edges are still coloured red, green or blue, edge operators have their Pauli bases defined according to their orientations within each T junction, i.e., horizontal edges are of $Z$ type while the $X$ and $Y$ checks are respectively those edges which are 90o clockwise and counter-clockwise from the horizontal edges (see Appendix A.1). All plaquettes have the same stabiliser operator, i.e., the XYZ2 operator. While the logical operators of the XYZ2 code have the same qubit supports as those of the P6 code, the qubit support bases of the XYZ2 logical operators are not uniform throughout, but involve $X$, $Y$, and $Z$ Paulis [6]. We note that even though these honeycomb code variants have been studied by several works in the literature (e,g., Refs. [6, 7, 8, 9]), there have not been any studies comparing the performance of these two codes under biased noise.
II.3 CSS Floquet codes
@@
Having discussed the honeycomb codes, we now give a brief review of another type of Floquet code, the CSS Floquet code which was proposed in Refs. [10, 11]. The CSS Floquet code is defined on the same honeycomb lattice. We show the code’s measurement cycle in Fig. 1(a) . We measure operators defined on edges using the $r\rightarrow g\rightarrow b$ cycle, but alternate between measuring $XX$ and $ZZ$ operators on these edges. Hence, this code has a period-6 measurement cycle, and its logical operator evolutions also have period-6. Even though a honeycomb code’s measurement schedule has only a period of 3, for consistency, we will define 1 QEC round to be 6 measurement subrounds for all codes studied in this paper. The check operators, stabilisers and logicals all are either $X$- or $Z$-type and, for this reason, the code is similar to static CSS codes.
@@
Unlike the honeycomb codes, the CSS Floquet code has no persistent stabiliser operators. Instead, each ISG contains $X$-type and $Z$-type stabiliser operators defined on two of the three colours of plaquette and either an $X$-type or a $Z$-type operator defined on the other colour. For instance, after measuring red-$X$ checks, the ISG contains $X^{\otimes 6}$ and $Z^{\otimes 6}$ blue and green plaquette operators, but only $X^{\otimes 6}$ red plaquette operators, since $Z^{\otimes 6}$ red plaquette operators anti-commute with the red-$X$ checks. See Appendix A for more details.
Plaquette operators are measured in a single measurement subround, and detectors are formed from the consecutive measurements of these plaquette operators, whenever their results can be compared. Plaquette operators anti-commute with check measurements at certain measurement subrounds and hence, we cannot necessarily form a detector at a plaquette every time a plaquette operator is measured (see Appendix A.2).
@@
The CSS Floquet code is naturally suited to minimum-weight perfect matching (MWPM) decoding, since single-qubit ($X$ or $Z$) Pauli errors and measurement errors all lead to graph-like syndromes [11]: they each trigger a pair of detectors. There are two decoding graphs formed from the $Z$-type and $X$-type detectors. Only $Y$ errors form hyperedges that need to be decomposed into edges in the two detector graphs.
@@
Figure 2: Persistent symmetry of the X3Z3 Floquet code leads to improved performance under $Z$-biased noise. (a) An example of such a symmetry, along with a Pauli-$Z$ error, is shown throughout the measurement cycle for the Floquet code. Arrows indicate the type of check measurements ($c\text{A}$ or $c$B for some colour $c$) performed in each subround. A-type plaquette operators that form part of the ISG at that subround are highlighted in darker colours. The product of these operators gives rise to a persistent symmetry under infinitely $Z$-biased noise. Anyons associated with a single Pauli-$Z$ error are shown at each subround: these always appear in pairs along the vertical strip in which the $Z$ error is contained. (b) An example of a symmetry after $r$A checks are measured. Since the product of red and blue A-type plaquettes along this vertical strip commutes with the noise, the syndromes must appear in pairs along this vertical strip.
@@
Figure 3: The A-type detector graph of the X3Z3 Floquet code under infinitely biased noise. Only a single vertical strip [e.g., that supporting the symmetry] is shown in Fig. 2(b). The illustrated part of the graph is disconnected from that of any neighbouring strip. The A-type measurement subrounds are shown with arrows on the left; the B-type subrounds are not shown, since they do not influence the detectors in this graph. Nodes (representing detectors) are placed at the position of the corresponding plaquette and at the subround in which it is measured to form a detector. For example, the bottom layer corresponds to blue A-check measurements, at which time red A-type plaquettes are measured.

### v2

@@ Abstract
We propose the X3Z3 Floquet code, a type of dynamical code with improved performance under biased noise compared to other Floquet codes. The enhanced performance is attributed to a simplified decoding problem resulting from a persistent symmetry under infinitely biased noise, which surprisingly exists in a code without constant stabilisers. Even if such a symmetry is allowed, we prove that a general dynamical code with two-qubit parity measurements cannot admit one-dimensional decoding graphs, a key feature resulting in the high performance of bias-tailored stabiliser codes. Despite this limitation, we demonstrate through our comprehensive numerical simulations that the symmetry of the X3Z3 Floquet code renders its performance under biased noise far better than several leading Floquet code candidates. Furthermore, to maintain high-performance implementation in hardwares without native two-qubit parity measurements, we introduce ancilla-assisted bias-preserving parity measurement circuits. Our work establishes the X3Z3 code as a prime quantum error-correcting code candidate, particularly for devices with reduced connectivity, such as the honeycomb and heavy-hexagonal architectures.
@@ I Introduction
Quantum error correction (QEC) [1, 2, 3, 4] should be understood as occurring both in space and time [5]. Taking advantage of the temporal dimension, Floquet codes [6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20], or more generally dynamical codes [21, 22, 23], form a large class of error-correcting codes, which can achieve competitive fault-tolerant performance [6, 8, 9, 10, 13, 12, 11, 14, 16, 15, 17] while reducing the weights of check measurements performed for the error correction. Several of these codes [7, 9, 18, 13] also benefit from being defined on a lattice with sparser connectivity than that for the surface code [24, 25]: each qubit is only connected to three other qubits. In architectures where two-qubit parity check measurements are native, Floquet codes could achieve higher thresholds [10, 8, 17] and lower space-time overheads [8, 10] than the surface code. Without requiring additional connectivities, this code can be deformed around defective components caused by highly noisy qubits or gates [15, 26]. Moreover, dynamical codes allow for implementations of arbitrary Clifford, and even some non-Clifford gates, through low-weight parity check measurements [21]. In addition, dynamical measurement schedules can result in certain errors being self-corrected [20].
@@
Although Floquet codes (in particular, the honeycomb code) have been studied under various noise models [9, 27], there have not been any Floquet codes that are specifically tailored for an improved performance under biased noise. A biased noise model is one in which a specific type of error, for example, phase errors, occurs more frequently than other errors, such as bit flip errors. This biased noise is typical to most quantum platforms, for example, superconducting fluxonium qubits [28], bosonic “cat” qubits [29, 30], spin-optical [31, 17], neutral atoms [32], quantum-dot spin qubits [33, 34] and Majorana qubits [35, 36].
@@
For enhanced performance, quantum-error-correcting codes need to be designed such that they possess symmetries that can be utilised to simplify the error-syndrome decoding problem given the noise structure [37, 38, 39, 40, 41, 42, 43, 44, 45]. While there have been a number of proposals for bias-tailored static codes [38, 39, 41, 45, 46, 40, 47, 41, 44, 42, 48, 49, 43, 50, 51], designing Floquet codes for high performance under biased noise is still an open problem. Owing to the experimental relevance of biased noise and given the ease of implementation of Floquet codes which require only two-qubit parity measurements, it is therefore imperative to tailor Floquet codes for biased noise and study how the performance of such dynamical codes can be improved.
@@
In this paper, we present the X3Z3 bias-tailored Floquet code, a Clifford-deformed [45] version of the Calderbank–Shor–Steane (CSS) Floquet code [18, 13]. Despite not having a fixed stabiliser group (as static codes have), the X3Z3 Floquet code surprisingly still possesses a symmetry under infinitely biased noise that allows for a simplified decoding in biased noise regimes. We perform an in-depth study of this code, along with the CSS Floquet code [18, 13], and two types of honeycomb codes: one introduced by Hastings and Haah [6] and the other by Gidney et al. [8]. We simulate all codes under biased-noise models, and find that the X3Z3 code has the best performance. Using a matching decoder, we find that, as the noise changes from fully depolarising to pure dephasing, the X3Z3 Floquet code threshold increases from $1.13\%$ to $3.09\%$ under a code-capacity noise model and increases from $0.76\%$ to $1.08\%$ under a circuit-level noise model mimicking hardwares with noisy direct entangling measurements. Furthermore, we show that its sub-threshold performance is also substantially better under biased noise than other Floquet codes.
@@
Compared to its static counterparts, the X3Z3 Floquet code has an advantage that it can be realised using only two-qubit parity check measurements. This makes it particularly suitable for devices with constrained connectivities, such as the honeycomb and heavy-hexagonal lattice (currently IBM’s preferred superconducting-qubit architecture) [52, 53]. Moreover, we demonstrate that the two-qubit parity measurements of the Floquet code can be performed in a bias-preserving way even in hardware without direct entangling measurements, thus enabling high performance implementation in such devices.
@@
The paper is laid out as follows. In Sec. II, we review the basics of Floquet codes together with two widely studied examples: honeycomb and CSS Floquet codes. Readers who are already familiar with Floquet codes can skip directly to Sec. III where we discuss our X3Z3 Floquet code which is obtained by bias-tailoring the CSS Floquet code. Crucially, we show that there exists a persistent symmetry in the code’s error syndrome under pure dephasing noise that allows for a simplified decoding. Subsequently, we introduce ancilla-assisted bias-preserving parity measurement circuits that allow for high-performance code implementation in devices without native entangling measurements. In Sec. IV, we present our simulation methods and noise models. In Sec. V, we then provide our simulation results for all of the four Floquet codes studied. To explain the code performance, in Sec. VI we prove that a class of dynamical codes on the honeycomb lattice has high-connectivity decoding graphs. We also prove that dynamical codes with only two-qubit parity check measurements cannot have repetition-code-like decoding graphs, as is the case with static codes. Finally, we conclude and present future research directions in Sec. VII. In the Appendices, we give a more detailed review of the basics of honeycomb and CSS Floquet codes, and present the details of our noise models, numerical simulations, and theorems.
@@ II.1 Dynamical and Floquet codes
Figure 1: (a) CSS and (b,c) X3Z3 Floquet codes. (a,b) Left: Hexagonal lattice with qubits sitting on vertices and opposite boundaries identified. Plaquettes and edges are assigned one of three colours [red, green, blue; examples are shown in (a)] and one of two types (X- or Z-type for CSS and A- or B-type for X3Z3). (a) Right: CSS Floquet code measurement schedule. (b) Left: The X3Z3 Floquet code is obtained from the CSS code via Hadamard gates applied to shaded domains’ qubits. Right: Plaquette and edge (check) operators are A- or B-type, depending on their support basis on the shaded/unshaded domains. (c) The X3Z3 Floquet code measurement schedule. Arrows indicate the type and colour of the edge operator measured at each step, where the edges just measured (members of the ISG) are highlighted in the lattice. Uncoloured plaquettes host only a single type of stabiliser, either A or B, indicated by the letters in the plaquettes, while coloured plaquettes host both A- and B-type stabilisers in the ISG. One set of anti-commuting logical operators is shown by yellow and brown strings, where their $X$ ($Z$) bases on the qubit supports are highlighted in red (blue). The other set (not shown) is similar to the set shown but offset by three measurement subrounds and with an $X\leftrightarrow Z$ interchange of the qubit supports’ bases.
@@
We begin by defining dynamical and Floquet codes. Here we consider the Floquet codes to be defined on the lattice of a two-dimensional colour code, which is trivalent and three-colourable. A trivalent lattice has each vertex incident to three edges, and a three-colourable lattice has every face assigned one of three colours in such a way that there are no two adjacent faces of the same colour. Throughout this paper, we will use the honeycomb lattice as an example of such a lattice (see Fig. 1).
@@
We define qubits as residing on the vertices of the lattice and error-correction will be performed by measuring two-qubit operators defined on edges of the lattice (i.e., acting on qubits incident to a given edge). Each edge is coloured the same as the plaquettes it connects [see e.g., the coloured edges in the middle right of the lattice in Fig. 1(a)]. We perform error-correction by measuring edge operators in a particular sequence. After any given subround of edge measurements, the system will be in the $+1$-eigenstate of the operators in an “instantaneous stabiliser group” (ISG), which will change at each time step. The ISG at time step $t+1$ is defined as
@@
$\begin{split}\text{ISG}_{t+1}=\langle S,\pm M\;|\;M\in\mathcal{M}_{t+1},\;S\in\text{ISG}_{t},\\ \text{such that }[S,M]=0,\;\forall M\in\mathcal{M}_{t+1}\rangle.\end{split}$ (1)
@@
In the above, $\mathcal{M}_{t+1}$ is the set of “check” measurements performed at time step $t+1$. The ISGt+1 also includes “plaquette stabiliser operators” $S\in\text{ISG}_{t}$ which commute with all $M\in\mathcal{M}_{t+1}$. The check measurement operators are chosen in such a way that those check operators at time $t$ that have overlapping qubit supports with check operators at $t+1$ anti-commute. For Floquet codes, the measurement sequence is periodic, such that $\mathcal{M}_{t+T}=\mathcal{M}_{t}$ for some integer $T$. For such a code, we will be performing quantum memory experiments with $mT$ time steps, for some integer $m$. We will refer to $m$ as the number of “QEC rounds” in the experiment, while we will refer to $mT$ as the number of “measurement subrounds” in the experiment.
@@
The logical operators at time $t$ are given by $\mathcal{C}(\mathrm{ISG}_{t})/\text{ISG}_{t}$, where $\mathcal{C}(\text{ISG}_{t})$ is the centraliser of ISGt, i.e., the group of Pauli operators commuting with all $S\in\text{ISG}_{t}$. A (potentially trivial) logical operator “representative” is some member of $\mathcal{C}(\text{ISG}_{t})$. Each nontrivial logical operator representative for the codes considered will be a string-like Pauli operator at each time step [see Fig. 1(c)]. To avoid anti-commuting with the next-subround edge measurements, certain check measurement results along a logical operator’s path have to multiplied into that logical operator. Hence, the logical operators will evolve from one time step to the next.
@@
We can detect errors if we can find sets of measurements, called detectors, that always multiply to the value $+1$ in the absence of noise, thus registering no error. Over some number of QEC rounds we will have extracted several detector outcomes. A detector (or decoding) hypergraph is formed by first defining a node for each (independent) detector in the code’s history. Subsequently, for each potential fault (e.g., Pauli or measurement errors) that might have occured, a (hyper)edge is drawn between the detectors whose signs are flipped by this fault. Each (hyper)edge is assigned a weight based on the probability of the corresponding error occurring [54]. The codes we will be examining are amenable to minimum-weight perfect matching decoding [55], upon decomposing hyperedges into edges. Given a “syndrome” (a set of detectors whose measurements return $-1$ rather than $+1$), the decoder attempts to pair up the triggered detectors to determine a highest-probability correction operation. The decoder succeeds if the error combined with the correction is a trivial logical operator.
@@ II.2 Honeycomb codes
We define one stabiliser operator for each plaquette, where red, green, and blue plaquettes host the $X^{\otimes 6}$, $Y^{\otimes 6}$, and $Z^{\otimes 6}$ operators, respectively. These plaquette operators are inferred from edge measurements in two consecutive subrounds. Detectors are formed from consecutive plaquette operator measurements.
@@
The second honeycomb code variant, which we call the XYZ2 honeycomb code, is due to Hastings and Haah [6]. It differs from the P6 code by single-qubit Clifford rotations acting on the qubits. While the XYZ2 code’s edges are still coloured the same as the P6 edges, the XYZ2 code’s edge operators have their Pauli bases defined according to their orientations within each T junction, i.e., horizontal edges are of $Z$ type while the $X$ and $Y$ checks are respectively those edges which are 90 clockwise and counter-clockwise rotated from the horizontal edges (see Appendix A.1). All plaquettes have the same stabiliser operator, i.e., the XYZ2 operator. While the logical operators of the XYZ2 code have the same qubit supports as those of the P6 code, the qubit support bases of the XYZ2 logical operators are not uniform throughout, but involve $X$, $Y$, and $Z$ Paulis [6]. We note that even though these honeycomb code variants have been studied by several works in the literature (e,g., Refs. [6, 8, 7, 9]), there have not been any studies comparing the performance of these two codes under biased noise.
II.3 CSS Floquet code
@@
Having discussed the honeycomb codes, we now give a brief review of another type of Floquet code, the CSS Floquet code, which was first presented in Refs. [18, 13]. We refer the readers to Appendix A.2 for more details on the code. The CSS Floquet code is defined on the same honeycomb lattice. We show the code’s measurement cycle in Fig. 1(a). We measure operators defined on edges using the $r\rightarrow g\rightarrow b$ cycle, but alternate between measuring $XX$ and $ZZ$ operators on these edges. Hence, this code has a period-6 measurement cycle, and its logical operator evolutions also have period-6 (see Appendix A.2 for an illustration). Even though a honeycomb code’s measurement schedule has only a period of 3, for consistency, we will define 1 QEC round to be 6 measurement subrounds for all codes studied in this paper. The check operators, stabilisers and logicals all are either $X$- or $Z$-type and, for this reason, the code is called the CSS Floquet code.
@@
Unlike the honeycomb codes, the CSS Floquet code has no persistent stabiliser operators. Instead, each plaquette stabiliser is repeatedly reinitialised and measured out, with subsequent measurement values compared to form a detector. In contrast to the honeycomb codes where the values of the plaquette stabilisers are inferred from measurements of check operators in two consecutive subrounds, here the plaquette operator measurements take only a single subround [13, 18]. Because of these single-step measurements, at each measurement subround, there is always one type of plaquette operator that anti-commutes with the check measurements, and hence their values are undetermined (see Appendix A.2). These “inactive” plaquettes, however, will be reinitialized at the next measurement subround. As a result, at each time step, the ISG contains $X$-type and $Z$-type stabiliser operators defined on two of the three colours of plaquette and either an $X$-type or a $Z$-type operator defined on the other colour. For instance, after measuring the red-$X$ checks, the ISG contains $X^{\otimes 6}$ and $Z^{\otimes 6}$ blue and green plaquette operators, but only $X^{\otimes 6}$ red plaquette operators, since $Z^{\otimes 6}$ red plaquette operators anti-commute with the red-$X$ checks (see Appendix A.2 for more details).
@@
The CSS Floquet code is naturally suited to a minimum-weight perfect matching (MWPM) decoding, since single-qubit ($X$ or $Z$) Pauli and measurement errors all lead to graph-like syndromes [13]: they each trigger a pair of detectors. There are two decoding graphs formed from the $Z$-type and $X$-type detectors. Only $Y$ errors form hyperedges that need to be decomposed into edges in the two detector graphs.
@@
Figure 2: Persistent symmetry of the X3Z3 Floquet code leads to improved performance under $Z$-biased noise. (a) An example of such a symmetry along a single unshaded strip of the lattice, together with a Pauli-$Z$ error, is shown throughout the measurement cycle. Arrows indicate the type of check ($c\text{A}$ or $c$B for some colour $c$) measured at each subround. A-type plaquette operators whose product forms the symmetry at that subround are highlighted in darker colours. (b) An example of a symmetry after $r$A checks are measured is the product of red and blue A-type plaquettes shown. Since this product is the identity along the vertical unshaded strip and commutes with $Z$ errors on all qubits, the syndromes/anyons [represented by coloured circles in (a)] must appear in pairs along the vertical strips.
@@
Figure 3: The A-type detector graph of the X3Z3 Floquet code under pure dephasing noise. Only a single vertical strip [e.g., that supporting the symmetry in Fig. 2(b)] is shown. The illustrated part of the graph is disconnected from that of any neighbouring strips. Arrows represent only the A-type measurements. The B-type subrounds are not shown, since they do not influence the detectors in this graph. Each node (representing a detector) is placed in the center of a plaquette and at the subround in which the corresponding detector is formed. For example, the bottom layer corresponds to the blue A-check measurements, at which time the red A-type plaquettes are measured.